OpenStudy (ifrah123):

show that 18! is congruent to -1 in modulo 437 ?

OpenStudy (latinc):

-1 mod 437 I believe equals 436; 436/2 = 218

OpenStudy (latinc):

idk how they are congruent

OpenStudy (latinc):

is there more to your question

OpenStudy (ifrah123):

it is solve through wilson's theorem which is (an integer p is a prime if and only if (p-1)! is congruent to -1 in modulo p )

jimthompson5910 (jim_thompson5910):

I'm noticing that 23*19 = 437 but I'm not sure how to fit it in

OpenStudy (latinc):

do you have the original question to post

ganeshie8 (ganeshie8):

23 and 19 both divide 218! Since they are primes, their product 23*19 also divides 218!

jimthompson5910 (jim_thompson5910):

wouldn't it be like this @ganeshie8 ? 218! = 218*217*...*24*23*22*21*20*19*18...*3*2*1 218! = 218*217*...*24*(23*19)*22*21*20*18...*3*2*1 218! = 218*217*...*24*(437)*22*21*20*18...*3*2*1 218! = 437*218*217*...*24*22*21*20*18...*3*2*1 but since 437 is a factor, this means 218! = 0 (mod 437) so I'm not sure if I messed up somewhere or if the original statement `218! = -1 (mod 437)` is incorrect?

ganeshie8 (ganeshie8):

Exactly The given congruence is false

OpenStudy (kainui):

Maybe they accidentally thought 437 was prime haha

jimthompson5910 (jim_thompson5910):

good point @Kainui I thought it was too until I used a prime number checker

OpenStudy (ifrah123):

sorry guys its 18! not 218!

jimthompson5910 (jim_thompson5910):

Here is one way to do it (see the attached text document). It's probably one of the longer ways possible. If not the longest route. There is probably a much more clever way to do this. I can't think of it right now.

ganeshie8 (ganeshie8):

Wilson gives us 18! = -1 (mod 19) 22! = -1 (mod 23)

ganeshie8 (ganeshie8):

Notice that 22! = 18!*19*20*21*22 = 18!*(-4)(-3)(-2)(-1) = 18!*24 = 18! (mod 23)

ganeshie8 (ganeshie8):

That means we have 18! = -1 (mod 19) 18! = -1 (mod 23) Since 19 and 23 are primes, 19*23 also divides 18!+1